LUCI1/AO Observing

Summary

Pupil shift on offset data sets

Offset Tracking number Pupil Offset Pyramid Movie
45" 20170119_050326 0.70 pupil_shift_45.mov
-60" 20170119_051132 -0.98 pupil_shift_m60.mov
-20" 20170119_053438 -0.48 pupil_offset_m20.mov
60" 20170119_053633 1.11 pupil_shift_60.mov

Details

00:30 SX WFS power on

00:47 collected 100 "Acquire HO Dark"
  • 20170119_004558
  • Looks like no extra noise in one quadrant of CCD39

04:30 Open

04:43 AO Closed loop bin 1
  • Optimize gain failed due to safe skip. Loop stayed closed

04:54 Close loop in intervention mode.
  • Skip optimize gain

04:55:46 offset, 45, 0
  • Loop paused and then "offsetXY - Command Interrupted"
  • Don't know why

05:03 closed loop
  • 45" offset
  • pupil_shift_45.mov
  • 20170119_050326

05:11 offset -60
  • 051132
  • pupil_shift_m60.mov

05:18 offset 60
  • out of field
  • data set not useful

05:24 Closed Loop
  • offset 30
  • 052429
  • resumed failed, 1.2 pixel diff in X
  • pupil_offset_30_2.mov

05:30 Use AOS Command GUI To close the loop because IIF thinks loop is closed
  • failed because of previous offset, no star

05:34 AO closed loop
  • offset -20
  • 053438
  • pupil_offset_m20.mov

05:36 Closed loop
  • offset 60
  • 053633
  • pupil_shift_60.mov

05:45 Failure during AcquireRefAO
  • sent 13.3, bin 2
  • found 15.2, bin 3
  • after changing binning on CCD39 then nothing
  • finally timed out after 5 minutes with "unknown exception from arbitrator" 05:52:02

05:57 waiting for PresetAO to timeout
  • WFS did not receive request
  • Unknown exception from arbitrator
  • stop start aoarb and adsecarb and wfsarb
  • AdSec Control GUI says AORunning
  • click StopAO, but nothing happens. wait wait wait
  • IDL process is stuck??

06:13
  • tried to stop start wfsarb
  • wfs msgrtdb died
  • had to w_stop/start
  • when power on get Communication error to Slope BCU
  • power off and on
  • OK

06:28 Unknown exception from arbitrator
  • PresetAO complete on AO Supervisor but AOS gets exception
  • tried from Command GUI and same problem

06:31 Steve stop/start AOS left

06:32 send LUCI preset, but shell was ripped due to AOS restart

06:45 BaysideX stage not homed.
  • homed it

06:55 PresetAO successful. Communication to AO Supervisor is now working fine

-- %USERSIG{DougMiller - 2017-01-19}%
Topic revision: r5 - 25 Jan 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback